### v2

# Discrete dark matter with light Dirac neutrinos

Abstract
We propose a new realisation of light Dirac neutrino mass and dark matter (DM) within the framework of a non-Abelian discrete flavour symmetry based on $A_{4}$ group. In addition to $A_{4}$, we also consider a $Z_{2}$ and an unbroken global lepton number symmetry $U(1)_{L}$ to keep unwanted terms away while guaranteeing the Dirac nature of light neutrinos. The field content, their transformations and flavon vacuum alignments are chosen in such a way that the type-I Dirac seesaw generates only one light Dirac neutrino mass while the other two masses arise from scotogenic contributions at one-loop. This leads to the Dirac scoto-seesaw framework, a generalisation of the widely studied scoto-seesaw model to Dirac neutrinos. The symmetry breaking of $A_{4}$ leaves a remnant $\mathcal{Z}_{2}$ symmetry responsible for stabilising DM. Dirac nature of light neutrinos introduces additional relativistic degrees of freedom $\Delta N_{\rm eff}$ within reach of cosmic microwave background experiments.
I Introduction
Origin of light neutrino mass and mixing ParticleDataGroup:2020ssz has been one of the longstanding puzzles in particle physics. While neutrinos remain massless within the framework of the standard model (SM) of particle physics, several beyond standard model (BSM) frameworks have been proposed to explain non-zero neutrino mass and mixing. Canonical seesaw frameworks Minkowski:1977sc ; GellMann:1980vs ; Mohapatra:1979ia ; Sawada:1979dis ; Yanagida:1980xy ; Schechter:1980gr ; Mohapatra:1980yp ; Lazarides:1980nt ; Wetterich:1981bx ; Schechter:1981cv ; Foot:1988aq involve inclusion of heavy fields which couple to the SM leptons. These seesaw mechanisms and several of their descendants predict Majorana light neutrino masses which violate lepton number by two units. However, there has been no confirmation about the Majorana nature of light neutrinos with experiments searching for neutrinoless double beta decay (NDBD) continuing to report null results. While null results at NDBD experiments do not necessarily imply Dirac nature of light neutrinos, there has been growing interest in Dirac neutrino models of late. After a few initial attempts in this direction Roncadelli:1983ty ; Roy:1983be ; Babu:1988yq ; Peltoniemi:1992ss , several new proposals to realise sub-eV scale Dirac neutrino masses Chulia:2016ngi ; Aranda:2013gga ; Chen:2015jta ; Ma:2015mjd ; Reig:2016ewy ; Wang:2016lve ; Wang:2017mcy ; Wang:2006jy ; Gabriel:2006ns ; Davidson:2009ha ; Davidson:2010sf ; Bonilla:2016zef ; Farzan:2012sa ; Bonilla:2016diq ; Ma:2016mwh ; Ma:2017kgb ; Borah:2016lrl ; Borah:2016zbd ; Borah:2016hqn ; Borah:2017leo ; CentellesChulia:2017koy ; Bonilla:2017ekt ; Memenga:2013vc ; Borah:2017dmk ; CentellesChulia:2018gwr ; CentellesChulia:2018bkz ; Han:2018zcn ; Borah:2018gjk ; Borah:2018nvu ; Borah:2019bdi ; CentellesChulia:2019xky ; Jana:2019mgj ; Nanda:2019nqy ; Guo:2020qin ; Bernal:2021ezl ; Borah:2022obi ; Li:2022chc ; Dey:2024ctx ; Singh:2024imk have appeared recently. Most of these works, in addition to the inclusion of new fields, also consider new discrete or continuous symmetries in order to forbid unwanted terms in the Lagrangian. In the spirit of Dirac seesaw for light neutrinos, such unwanted terms include the bare mass term of Majorana type for singlet right handed neutrinos $(\nu_{R})$ and direct coupling of $\nu_{R}$ with the SM lepton and Higgs doublets.
Another observed phenomena which can not be explained in the SM is the presence of dark matter (DM) in the Universe. Its presence is supported by astrophysical observations at different scales Zwicky:1933gu ; Rubin:1970zza ; Clowe:2006eq together with cosmological experiments like PLANCK, WMAP predicting around 26.8% of the present Universe to be made up of DM ParticleDataGroup:2020ssz ; Planck:2018vyg . While weakly interacting massive particle (WIMP) has been the most widely studied particle DM framework, null results at direct detection experiments LUX-ZEPLIN:2022qhg have also led to growing interest in alternative scenarios where DM is more feebly coupled to the SM, like the feebly interacting massive particle (FIMP). Recent reviews of WIMP and FIMP can be found in Arcadi:2017kky and Bernal:2017kxu respectively.
Motivated by this, we propose a new flavour symmetric origin of light Dirac neutrino mass and dark matter in this work. We consider the popular non-Abelian discrete flavour symmetry group of $A_{4}$ augmented with an additional discrete $Z_{2}$ and global lepton number $U(1)_{L}$ symmetries. The $A_{4}$ flavour symmetry gets spontaneously broken by flavon fields while generating light Dirac neutrino mass and mixing. For a review of non-Abelian discrete flavour symmetries and their consequences, see Ref. Altarelli:2010gt ; King:2013eh ; Petcov:2017ggy ; Chauhan:2023faf ; Ding:2024ozt and references therein. Unlike popular DM scenarios with additional stabilising symmetries, here we have stable DM by virtue of a remnant $\mathcal{Z}_{2}$ symmetry naturally emerging from $A_{4}$ breaking. Similar scenario for Majorana light neutrino, coined as discrete dark matter was proposed in Hirsch:2010ru and subsequently explored within different contexts Meloni:2010sk ; Boucenna:2011tj ; Adulpravitchai:2011ei ; Eby:2011qa ; Boucenna:2012qb ; Hamada:2014xha ; Lamprea:2016egz ; DeLaVega:2018bkp ; Bonilla:2023pna ; Kumar:2024zfb . We propose a realisation of this discrete dark matter paradigm with light Dirac neutrinos in a minimal $A_{4}$ flavour symmetric setup. We show that neutrino oscillation data can be satisfied after taking tree-level and one-loop contributions to light Dirac neutrino masses. Interestingly, this also leads to the generalisation of the popular Scoto-Seesaw scenario Rojas:2018wym to Dirac neutrinos, and can be termed as Dirac Scoto-Seesaw. While the model predicts either scalar or fermion DM with the desired relic abundance, light Dirac neutrinos lead to observable $\Delta N_{\rm eff}$ at future cosmic microwave background (CMB) experiments like CMB Stage IV (CMB-S4) Abazajian:2019eic , SPT-3G SPT-3G:2014dbx , Simons Observatory SimonsObservatory:2018koc and CMB-HD CMB-HD:2022bsz .
This paper is organised as follows. In section II, we discuss our model and the origin of light Dirac neutrino mass with stable dark matter. In section III, we discuss the numerical analysis related to neutrino mass and mixing followed by discussion of dark matter, $\Delta N_{\rm eff}$ and other detection prospects in section IV, V, VI respectively. Finally, we conclude in section VII.
II The Model
In this section, we discuss the relevant field content of the model, the field transformations under the chosen symmetries and the particle spectrum. Table 1 summarises the relevant field content and the corresponding quantum numbers. The SM lepton doublets and right chiral part of Dirac neutrino $(\nu_{R})$ transform as $A_{4}$ singlets while heavy vector-like SM gauge singlet fermions $N_{L,R}$ transform as $A_{4}$ triplets. In addition to the SM Higgs $H$ transforming as $A_{4}$ singlet, additional scalar doublet $\eta$ and scalar singlet $\chi$ both transforming as $A_{4}$ triplets are included. The additional $Z_{2}$ symmetry prevents direct coupling of SM lepton doublet, SM Higgs doublet and $\nu_{R}$. On the other hand, a global unbroken lepton number symmetry $U(1)_{L}$ prevent Majorana mass terms guaranteeing the Dirac nature of light neutrinos. The Yukawa Lagrangian for charged lepton can be written as
$-\mathcal{L}_{\rm CL}\supset Y_{e}\overline{\ell_{e}}He_{R}+Y_{\mu}\overline{\ell_{\mu}}H\mu_{R}+Y_{\tau}\overline{\ell_{\tau}}H\tau_{R}+{\rm h.c.}$ (1)
Fields $\ell_{e}$ $\ell_{\mu}$ $\ell_{\tau}$ $e_{R}$ $\mu_{R}$ $\tau_{R}$ $N_{L,R}$ $\nu_{R_{e,\mu,\tau}}$ $\eta$ $\chi$ $H$
$SU(2)$ 2 2 2 1 1 1 1 1 2 1 2
$A_{4}$ $1$ $1^{\prime\prime}$ $1^{\prime}$ $1$ $1^{\prime\prime}$ $1^{\prime}$ $3$ $1$ , $1^{\prime}$ ,$1^{\prime\prime}$ $3$ $3$ $1$
$Z_{2}$ $1$ 1 1 $1$ $1$ $1$ $1$ $-1$ $1$ $-1$ $1$
$U(1)_{L}$ 1 1 1 1 1 1 1 1 0 0 0
Table 1: Field contents and transformation under the symmetries of our model.
On the other hand, the neutral fermion Lagrangian is given by
$-\mathcal{L}_{\nu}$ $\supset$ ${y_{1}}\bar{\ell_{e}}(N_{R}\tilde{\eta})_{1}+{y_{2}}\bar{\ell_{\mu}}(N_{R}\tilde{\eta})_{1^{\prime\prime}}+{y_{3}}\bar{\ell_{\tau}}(N_{R}\tilde{\eta})_{1^{\prime}}$
$+{M_{N}(\bar{N_{L}}N_{R})_{1}}+y^{\prime}_{1}(\bar{N}_{L}\chi)_{1}\nu_{Re}+y^{\prime}_{2}(\bar{N}_{L}\chi)_{1^{\prime\prime}}\nu_{R\mu}+y^{\prime}_{3}(\bar{N}_{L}\chi)_{1^{\prime}}\nu_{R\tau}+{\rm h.c.}$
$=$ ${y_{1}}(\bar{\ell_{e}}N_{R_{1}}\tilde{\eta_{1}}+\bar{\ell_{e}}N_{R_{2}}\tilde{\eta_{2}}+\bar{\ell_{e}}N_{R_{3}}\tilde{\eta_{3}})+{y_{2}}(\bar{\ell_{\mu}}N_{R_{1}}\tilde{\eta_{1}}+\omega\bar{\ell_{\mu}}N_{R_{2}}\tilde{\eta_{2}}+\omega^{2}\bar{\ell_{\mu}}N_{R_{3}}\tilde{\eta_{3}})$
$+{y_{3}}(\bar{\ell_{\tau}}N_{R_{1}}\tilde{\eta_{1}}+\omega^{2}\bar{\ell_{\tau}}N_{R_{2}}\tilde{\eta_{2}}+\omega\bar{\ell_{\tau}}N_{R_{3}}\tilde{\eta_{3}})+M_{N}\bar{N}_{L_{1}}N_{R_{1}}+M_{N}\bar{N}_{L_{2}}N_{R_{2}}+M_{N}\bar{N}_{L_{3}}N_{R_{3}}$
$+y^{\prime}_{1}(\bar{N}_{L_{1}}\chi_{1}+\bar{N}_{L_{2}}\chi_{2}+\bar{N}_{L_{3}}\chi_{3})\nu_{Re}+y^{\prime}_{2}(\bar{N}_{L_{1}}\chi_{1}+\omega\bar{N}_{L_{2}}\chi_{2}+\omega^{2}\bar{N}_{L_{3}}\chi_{3})\nu_{R\mu}$
$+y^{\prime}_{3}(\bar{N}_{L_{1}}\chi_{1}+\omega^{2}\bar{N}_{L_{2}}\chi_{2}+\omega\bar{N}_{L_{3}}\chi_{3})\nu_{R\tau}+{\rm h.c.}$ (3)
where the $A_{4}$ product rules, given in Appendix A, have been used to write the interactions in component form. The scalar potential of the model is given in Appendix B. The Yukawa Lagrangian for charged leptons given by Eq. (1) gives rise to a diagonal charged lepton mass matrix.
Figure 1: Neutrino mass generation at tree level and one-loop level.
At tree level the neutral fermion mass matrices in the ($\nu_{L},N_{R}$), ($N_{L},\nu_{R}$), $(N_{L},N_{R})$ basis can be written as
$m_{D}=\frac{v_{\eta}}{\sqrt{2}}\begin{pmatrix}y_{1}&0&0\\ y_{2}&0&0\\ y_{3}&0&0\end{pmatrix},m^{\prime}_{D}=\frac{v_{\chi}}{\sqrt{2}}\begin{pmatrix}y^{\prime}_{1}&y^{\prime}_{2}&y^{\prime}_{3}\\ 0&0&0\\ 0&0&0\end{pmatrix},M_{R}=\begin{pmatrix}M_{N}&0&0\\ 0&M_{N}&0\\ 0&0&M_{N}\end{pmatrix};$ (4)
respectively, where the vacuum expectation values (VEVs) of the $A_{4}$ scalar triplets are considered to be $\langle\eta\rangle=(\langle\eta_{1}\rangle,\langle\eta_{2}\rangle,\langle\eta_{3}\rangle)=(\frac{v_{\eta}}{\sqrt{2}},0,0)$ and $\langle\chi\rangle=(\langle\chi_{1}\rangle,\langle\chi_{2}\rangle,\langle\chi_{3}\rangle)=(\frac{v_{\chi}}{\sqrt{2}},0,0)$. Denoting the VEV of the neutral component of the Higgs doublet $H$ as $v$, we have the constraint $\sqrt{v^{2}_{\eta}+v^{2}}=246$ GeV from the requirement of desired electroweak symmetry breaking (EWSB). Therefore, at three level, via type-I seesaw like scenario (shown in the left panel of Fig. 1) , the light Dirac neutrino mass matrix can be written as
$m_{\nu 0}=-m_{D}M^{-1}_{R}m^{\prime}_{D}=-\frac{v_{\eta}v_{\chi}}{2M_{N}}\begin{pmatrix}y_{1}y^{\prime}_{1}&y_{1}y^{\prime}_{2}&y_{1}y^{\prime}_{3}\\ y_{2}y^{\prime}_{1}&y_{2}y^{\prime}_{2}&y_{2}y^{\prime}_{3}\\ y_{3}y^{\prime}_{1}&y_{3}y^{\prime}_{2}&y_{3}y^{\prime}_{3}\end{pmatrix}.$ (5)
Clearly, the tree level neutrino mass matrix is a matrix of rank-1 predicting only one massive neutrino with mass $m_{0}=-\frac{v_{\eta}v_{\chi}}{2M_{N}}(y_{1}y^{\prime}_{1}+y_{2}y^{\prime}_{2}+y_{3}y^{\prime}_{3})$.
At one-loop, all the $A_{4}$ triplet scalars $\eta=(\eta_{1},\eta_{2},\eta_{3})$ and $\chi=(\chi_{1},\chi_{2},\chi_{3})$ irrespective of their VEVs can contribute to light neutrino mass, as shown in the right panel of Fig. 1. This contribution is similar to the Dirac scotogenic [^1] Scotogenic model was originally proposed in Tao:1996vb ; Ma:2006fn for Majorana neutrinos. model discussed in several earlier works Gu:2007ug ; Farzan:2012sa ; Borah:2016zbd ; Ma:2016mwh ; Borah:2017leo ; Wang:2017mcy ; Ma:2019iwj ; Ma:2019yfo ; Ma:2019coj ; Leite:2020wjl ; Guo:2020qin ; Bernal:2021ezl ; Chowdhury:2022jde ; Borah:2022phw ; Borah:2022enh . The one-loop contribution to light Dirac neutrino mass can be estimated as
$m^{\rm loop}_{\nu}=\sum^{3}_{j=1}Y_{kj}Y^{\prime}_{jl}M_{j}\mathcal{F}_{j},$ (6)
where
$\mathcal{F}_{j}=\sum_{k,l}\frac{\sin\theta_{kl}\cos\theta_{kl}}{32\pi^{2}}\left(\frac{m^{2}_{\xi_{k}}}{m^{2}_{\xi_{k}}-M^{2}_{j}}{\rm ln}\frac{m^{2}_{\xi_{k}}}{M^{2}_{j}}-\frac{m^{2}_{\xi_{l}}}{m^{2}_{\xi_{l}}-M^{2}_{j}}{\rm ln}\frac{m^{2}_{\xi_{l}}}{M^{2}_{j}}\right),$ (7)
with $m_{\xi_{k}},m_{\xi_{l}}$ being the mass eigenstates of the associated scalars, $\theta_{kl}$ being the mixing angle and $M_{j}=M_{N}$. Following Eq. (3), the Yukawa matrices $Y$, $Y^{\prime}$ involved in the scotogenic contribution above can be written as
$Y=\begin{pmatrix}y_{1}&y_{1}&y_{1}\\ y_{2}&\omega y_{2}&\omega^{2}y_{2}\\ y_{3}&\omega^{2}y_{3}&\omega y_{3}\end{pmatrix},~{}~{}Y^{\prime}=\begin{pmatrix}y^{\prime}_{1}&y^{\prime}_{2}&y^{\prime}_{3}\\ y^{\prime}_{1}&\omega y^{\prime}_{2}&\omega^{2}y^{\prime}_{3}\\ y^{\prime}_{1}&\omega^{2}y^{\prime}_{2}&\omega y^{\prime}_{3}\end{pmatrix}.$ (8)
Using these, the complete one-loop mass matrix for light Dirac neutrinos can be written as
$m^{\rm loop}_{\nu}$ $=$ $Y_{k1}Y^{\prime}_{1l}M_{N}\mathcal{F}_{1}+Y_{k2}Y^{\prime}_{2l}M_{N}\mathcal{F}_{2}+Y_{k3}Y^{\prime}_{3l}M_{N}\mathcal{F}_{3}$
$=$ $M_{N}\begin{pmatrix}y_{1}y^{\prime}_{1}(\mathcal{F}_{1}+\mathcal{F}_{2}+\mathcal{F}_{3})&y_{1}y^{\prime}_{2}(\mathcal{F}_{1}+\omega\mathcal{F}_{2}+\omega^{2}\mathcal{F}_{3})&y_{1}y^{\prime}_{3}(\mathcal{F}_{1}+\omega^{2}\mathcal{F}_{2}+\omega\mathcal{F}_{3})\\ y_{2}y^{\prime}_{1}(\mathcal{F}_{1}+\omega\mathcal{F}_{2}+\omega^{2}\mathcal{F}_{3})&y_{2}y^{\prime}_{2}(\mathcal{F}_{1}+\omega^{2}\mathcal{F}_{2}+\omega\mathcal{F}_{3})&y_{2}y^{\prime}_{3}(\mathcal{F}_{1}+\mathcal{F}_{2}+\mathcal{F}_{3})\\ y_{3}y^{\prime}_{1}(\mathcal{F}_{1}+\omega^{2}\mathcal{F}_{2}+\omega\mathcal{F}_{3})&y_{3}y^{\prime}_{2}(\mathcal{F}_{1}+\mathcal{F}_{2}+\mathcal{F}_{3})&y_{3}y^{\prime}_{3}(\mathcal{F}_{1}+\omega\mathcal{F}_{2}+\omega^{2}\mathcal{F}_{3})\end{pmatrix}$
Therefore, the light Dirac neutrino mass matrix, taking tree level and one-loop contributions, can be written as
$m_{\nu}$ $=$ $m_{\nu 0}+m^{\rm loop}_{\nu}=\begin{pmatrix}(m_{\nu})_{11}&(m_{\nu})_{12}&(m_{\nu})_{13}\\ (m_{\nu})_{21}&(m_{\nu})_{22}&(m_{\nu})_{23}\\ (m_{\nu})_{31}&(m_{\nu})_{32}&(m_{\nu})_{33}\end{pmatrix}$ (11)
with the details of the components $(m_{\nu})_{ij}$ being given in Appendix C. For simplicity, the Yukawa couplings appearing in the neutrino mass matrix are considered to be real. Thus, CP violation is introduced only through the complex parameter $\omega$, appearing here as a consequence of the considered $A_{4}$ discrete flavour symmetry. This mass matrix has rank 3 and predict three non-zero mass eigenvalues, in general. Also, it does not possess any specific symmetric structure and hence can be fitted to neutrino mixing data by appropriate choice of parameters, as we discuss below. Interestingly, in the limit $y_{2}=y_{3}$ and $y^{\prime}_{2}=y^{\prime}_{3}$, the effective mass matrix given in Eq. (11) yields $\mu-\tau$ reflection symmetric mass matrix Harrison:2002et predicting fixed values for $\theta_{23}(=\pi/4)$ and $\delta_{\rm CP}(=\pm\pi/2$), see Appendix C for details. Since type-I seesaw generates only one mass scale with the other being generated by the scotogenic contribution, our framework provides a generalisation of the scoto-seesaw mechanism Rojas:2018wym ; Mandal:2021yph ; Barreiros:2020gxu ; Barreiros:2022aqu ; Ganguly:2022qxj ; Ganguly:2023jml ; Kumar:2023moh to light Dirac neutrinos. Similar to earlier works on Majorana neutrinos, Dirac scoto-seesaw framework can explain the two mass scales in neutrino sector namely, atmospheric and solar mass scales, from tree level (type-I seesaw) and radiative (scotogenic) contributions respectively.
It should be noted that the above choice of VEV alignment leaves a remnant $\mathcal{Z}_{2}\subset A_{4}$ unbroken after spontaneous symmetry breaking. The following fields transform non-trivially under this remnant symmetry as
$\mathcal{Z}_{2}:\,\,\eta_{2,3}\rightarrow-\eta_{2,3};\,\,\,\,\chi_{2,3}\rightarrow-\chi_{2,3};\,\,\,\,N_{2,3}\rightarrow-N_{2,3}$ (12)
while all other fields are $\mathcal{Z}_{2}$-even. The lightest $\mathcal{Z}_{2}$-odd particle, therefore, is stable and can be a dark matter candidate, the details of which are discussed in section IV.
III Fit to Neutrino data
[…]
Figure 2: Allowed regions for $y_{i}$, $y^{\prime}_{i}$ for the $3\sigma$ ranges for neutrino oscillation data for $y_{i}\gg y^{\prime}_{i}$ (Case-A).
[…]
Figure 3: Correlation for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ for $y_{i}\gg y^{\prime}_{i}$ (Case-A).
data Esteban:2020cvm , the allowed ranges for $y_{i},y^{\prime}_{i}$ are given in $y^{\prime}_{1}-y_{1}$, $y^{\prime}_{2}-y_{2}$ and $y^{\prime}_{3}-y_{3}$ planes respectively, plotted in Fig. 2. In this benchmark scenario, $y^{\prime}_{i}$ are found to be in the range $6\times 10^{-11}\leq y^{\prime}_{1}\leq 10^{-10}$, $3\times 10^{-11}\leq y^{\prime}_{2,3}\leq 8.5\times 10^{-11}$ and corresponding ranges for $y_{i}$ are found to be $0.03\leq y_{1}\leq 0.05$, $0.052\leq y_{2}\leq 0.1$ and $0.05\leq y_{3}\leq 0.1$ respectively. To be in agreement with neutrino oscillation data, other parameters are restricted within the range $2.8\,\leq(M_{N}/{\rm TeV})\leq 5$, $20\leq(v_{\eta}/{\rm GeV})\leq 240$, $0.22\leq(v_{\chi}/{\rm TeV})\leq 1.8$ with the loop functions defined in Eq. (7) having the magnitude $\mathcal{F}_{i}\sim\mathcal{O}(10^{-4})$. With the estimation of the parameters appearing in the neutrino mass matrix, we are now in a position to show the correlation among neutrino oscillation parameters as shown in Fig. 3. In this figure, we have plotted the correlations for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ planes for Case-A and predicted ranges for $\delta_{CP}$, $\sum m_{i}$ and $m_{1}$ are found to be in the range $(0.43-1.6)$ radian, $(0.06-0.09)$ eV and $(8.4-9.7)$ meV respectively.
Case-B: We now consider the benchmark scenario, $y_{i}\ll y^{\prime}_{i}$. Using similar methodology, we can again constrain the Yukawa couplings $y_{i},y^{\prime}_{i}$ imposing the constraints on neutrino masses and mixing angles Esteban:2020cvm .
Figure 4: Allowed regions for $y_{i}$, $y^{\prime}_{i}$ for the $3\sigma$ ranges for neutrino oscillation data for $y_{i}\ll y^{\prime}_{i}$ (Case-B).
In Fig. 4 we have plotted the allowed values of the respective Yukawa couplings in $y^{\prime}_{1}-y_{1}$, $y^{\prime}_{2}-y_{2}$ and $y^{\prime}_{3}-y_{3}$ planes and $y_{i},y^{\prime}_{i}$ are found to be restricted within the range $1.4\times 10^{-11}\leq y_{1}\leq 2.5\times 10^{-11}$, $2\times 10^{-11}\leq y_{2}\leq 7\times 10^{-11}$, $3.4\times 10^{-11}\leq y_{3}\leq 5.5\times 10^{-11}$
Figure 5: Correlation for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$ and $\sum m_{i}-\sin^{2}\theta_{23}$ for $y_{i}\ll y^{\prime}_{i}$ (Case-B).
and $0.01\leq y^{\prime}_{1}\leq 0.1$, $0.03\leq y^{\prime}_{2}\leq 0.1$, $0.04\leq y^{\prime}_{3}\leq 0.1$ respectively. In order to satisfy the neutrino oscillation data, other parameters are restricted within the range $6\,\leq(M_{N}/{\rm TeV})\leq 10$, $20\leq(v_{\eta}/{\rm GeV})\leq 240$, $0.2\leq(v_{\chi}/{\rm TeV})\leq 1.8$. Corresponding loop functions $\mathcal{F}_{i}$, defined in Eq. (7), appearing in the Dirac scotogenic contribution are found to be in the range $5\times 10^{-4}\leq\mathcal{F}_{1}\leq 7\times 10^{-4}$, $3\times 10^{-4}\leq\mathcal{F}_{2}\leq 10^{-3}$, $8\times 10^{-4}\leq\mathcal{F}_{3}\leq 1.3\times 10^{-3}$. After estimating the parameters relevant for light neutrino mass matrix, we proceed to find some correlations among neutrino oscillation parameters as shown in Fig. 5. In this figure, we have plotted the correlations for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ planes for the scenario $y_{i}\ll y^{\prime}_{i}$ and predicted ranges for $\delta_{CP}$, $\sum m_{i}$ and $m_{1}$ are found to be in the range ±$\pm(0-1.6)$ radian, $(0.068-0.11)$ eV and $(8.4-10.5)$ meV respectively.
Case-C: Finally, we now consider the third benchmark scenario where $y_{i}\simeq y^{\prime}_{i}$. In this scenario, one can also satisfy the observed constraints on neutrino masses and mixing
Figure 6: Allowed regions for $y_{i}$, $y^{\prime}_{i}$ for the $3\sigma$ ranges for neutrino oscillation data for $y_{i}\simeq y^{\prime}_{i}$ (Case-C).
where the associated Yukawa couplings are found to be in the range $3\times 10^{-7}\leq y_{1}\leq 4.6\times 10^{-7}$, $6.9\times 10^{-7}\leq y_{2}\leq 7.9\times 10^{-7}$, $7.1\times 10^{-7}\leq y_{3}\leq 8.1\times 10^{-7}$ and $5.6\times 10^{-6}\leq y^{\prime}_{1}\leq 6.8\times 10^{-6}$, $2.9\times 10^{-6}\leq y^{\prime}_{2}\leq 5.5\times 10^{-6}$ and $3.6\times 10^{-6}\leq y^{\prime}_{3}\leq 4.75\times 10^{-6}$ respectively. These allowed ranges are plotted in $y^{\prime}_{i}-y_{i}$ planes as given in Fig. 6.
Figure 7: Correlation for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$ and $\sum m_{i}-\sin^{2}\theta_{23}$ for $y_{i}\simeq y^{\prime}_{i}$ (Case-C).
To fit the model with the neutrino oscillation data in this scenario, other parameters are restricted within the range $8\,\leq(M_{N}/{\rm TeV})\leq 16$, $35\leq(v_{\eta}/{\rm GeV})\leq 240$, $0.23\leq(v_{\chi}/{\rm TeV})\leq 1.8$. Similarly, loop functions $\mathcal{F}_{i}$ defined in Eq. (7), appearing in the Dirac scotogenic contribution are found to be in the range $10^{-3}\leq\mathcal{F}_{1}\leq 1.2\times 10^{-3}$, $4\times 10^{-4}\leq\mathcal{F}_{2}\leq 7\times 10^{-4}$, $10^{-4}\leq\mathcal{F}_{3}\leq 1.7\times 10^{-4}$. Similar to the other two cases, here also we show the correlation among neutrino oscillation parameters in Fig. 7. In this figure, we have plotted the correlations for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ planes for the current scenario and predicted ranges for $\delta_{CP}$, $\sum m_{i}$ and $m_{1}$ are found to be in the range $(1.2-1.58)$ and $-(1-1.58)$ radian, $(0.082-0.11)$ eV and $(8.8-11)$ meV respectively.
In the following sections, for compactness, we will consider these three benchmark scenarios for studies on scalar/fermionic WIMP dark matter, $\Delta N_{\rm eff}$ and LFV decays associated with this discrete dark matter model for Dirac neutrinos.
IV Dark matter
In this model, gauge singlet Dirac fermions $N_{2,3}$, gauge singlet scalars $\chi_{2,3}$ and $SU(2)$ doublet scalars $\eta_{2,3}$, remain odd under the residual $\mathcal{Z}_{2}$ symmetry such that the lightest among them becomes a viable dark matter candidate. When considering the fermionic dark matter scenario, $i.e.$, $M_{N}<M_{\eta}<M_{\chi}$, achieving the correct relic density becomes challenging because the mass $M_{N}$ is constrained to the multi-TeV range by the neutrino mass criteria. Fermion DM in scotogenic model was studied in several earlier works, see Ahriche:2017iar ; Borah:2018smz ; Mahanta:2019gfe for example. As can be seen from these earlier works, fermion singlet DM relic relies primarily on Yukawa portal annihilations and coannihilations. Due to the involvement of the same Yukawa couplings in both type-I and scotogenic contributions to neutrino mass, we can not have Yukawa coupling in $\geq\mathcal{O}(1)$ regime as discussed in the previous section. After including coannihilation among fermion DM $N$ and $\mathcal{Z}_{2}$-odd scalars, correct relic abundance is found only for Case-A with $M_{N}\in[3000,5000]$ GeV, considering the parameter space in agreement with neutrino data. However, for Case-B ($M_{N}\in[6000,10000]$ GeV) and Case-C ($M_{N}\in[11400,14000]$ GeV), where $M_{N}$ is pushed to even larger values, it is not possible to achieve the correct relic density for the fermionic dark matter via thermal freeze-out. This is because the large $M_{N}$ value and insufficient (co)annihilation rates always lead to an overabundant relic density. Thermal relic of $N$ in Case-A is obtained due to smaller $M_{N}$ and large co-annihilation rate with $SU(2)$ doublet scalars via large Yukawa portal ($y_{i}\gg y^{\prime}_{i}$) interactions. Since doublet scalars have more efficient annihilation rates, it is easier to obtain correct relic of $N$ in Case-A compared to Case-B where $N$ has sizeable Yukawa with scalars which are predominantly singlets. While Case-C offers the possibility of non-thermal or freeze-in fermion singlet DM, we find it to be overproduced as well leaving only Case-A as viable for fermion DM scenario.
Figure 8: Left panel: variation of scalar DM relic abundance with its mass. The blue solid line indicates the observed DM relic. Right panel: spin-independent DM-nucleon scattering as a function of the DM mass. The red colored points satisfy DM relic. The cyan shaded region is ruled out by direct detection constraints.
In the case of scalar DM, the correct relic density can be obtained in all the three cases discussed in the previous section due to the gauge and scalar portal annihilation and co-annihilation of the scalar DM while successfully explaining the neutrino data. Due to the possibility of additional portals namely gauge and scalar portals for annihilation, scalar DM relic can be obtained without relying on the Yukawa portal, unlike fermion singlet DM. With $M_{N}$ varied in the required range so as to satisfy the neutrino oscillation data, we vary the masses of $\mathcal{Z}_{2}$-odd scalars $\zeta_{i}$ (see Appendix B for the details of physical scalar masses and eigenstates) assuming two hierarchies $M_{\zeta_{i}}<M_{N}$ or $M_{\zeta_{1}}<M_{N}<M_{\zeta_{j}},(j\neq 1)$. We consider $\zeta_{1}$, the lightest $\mathcal{Z}_{2}$-odd scalar to be the DM. As the mass difference between $\zeta_{i}$’s are constrained in a specific range from the neutrino mass criteria, when the mass hierarchy is $M_{\zeta_{i}}<M_{N}$, it is possible to get DM masses in the sub-TeV range. However when we assume the hierarchy $M_{\zeta_{1}}<M_{N}<M_{\zeta_{j}}(j\neq 1)$, the DM mass always lies in the multi-TeV scale.
In the left panel of Fig. 8, we show the variation of scalar DM relic with its mass. As shown in several earlier works on scalar DM, the (co)annihilations mediated by electroweak gauge bosons and the SM Higgs are responsible for the dip for intermediate mass, leaving two distinct mass ranges satisfying correct relic. The right panel of the same figure shows the spin-independent DM-nucleon scattering cross-section (mediated via Higgs portal) as a function of scalar DM mass for the points which satisfy relic abundance criteria. We used the package MicrOMEGAs Alguero:2023zol to calculate the relic density and DM-nucleon scattering cross-section after generating the model files using LanHEP Semenov:2014rea . The cyan shaded region is ruled out by the most stringent direct search constraints from LZ-2022 LUX-ZEPLIN:2022qhg for heavier DM and CRESST-III CRESST:2019jnq and DarkSide-50 DarkSide:2018bpj for lighter DM. Clearly we can see that the viable DM mass range that satisfies correct relic density and direct detection rate ranges from around $500$ GeV to several TeV scale. Interestingly, future direct detection experiment Darwin DARWIN:2016hyl can probe a part of this high mass region up to $M_{\rm DM}\sim 3$ TeV.
V Observable $\Delta N_{\rm eff}$
The effective number of relativistic degrees of freedom ${\rm N}_{\rm eff}$ defined as
${\rm N}_{\rm eff}=\frac{8}{7}\left(\frac{11}{4}\right)^{4/3}\left(\frac{\rho_{R}-\rho_{\gamma}}{\rho_{\gamma}}\right)\,,$ (13)
where $\rho_{R},\rho_{\gamma}$ denote total radiation and photon densities respectively. Existing data from CMB experiments like PLANCK constrains such additional light species by putting limits on the effective degrees of freedom for neutrinos during the era of recombination ($z\sim 1100$) as Planck:2018vyg
${\rm N_{eff}=2.99^{+0.34}_{-0.33}}$ (14)
at $2\sigma$ or $95\%$ C.L. including baryon acoustic oscillation (BAO) data. At $1\sigma$ C.L. it becomes more stringent to ${\rm N}_{\rm eff}=2.99\pm 0.17$. Similar bound also exists from big bang nucleosynthesis (BBN), $3.42.3<{\rm N}_{\rm eff}<$ 3.4 at $95\%$ C.L. Cyburt:2015mya . Recent observations by the DESI collaboration report $\Delta N_{\rm eff}<0.4$ at $2\sigma$ DESI:2024mwx , which is weaker than the PLANCK bound given in Eq. (14). All these bounds are consistent with SM predictions ${\rm N^{SM}_{eff}}=3.045$ Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq . Future CMB experiment CMB-S4 is expected reach a much better sensitivity of $\Delta{\rm N}_{\rm eff}={\rm N}_{\rm eff}-{\rm N}^{\rm SM}_{\rm eff}=0.06$ Abazajian:2019eic , taking it closer to the SM prediction. Another future CMB experiment, specially, CMB-HD CMB-HD:2022bsz can probe $\Delta N_{\rm eff}$ up to $0.014$ at $1\sigma$. Dirac nature of light neutrinos introduces additional relativistic degrees of freedom in the form of right chiral neutrinos which can contribute to $\Delta{\rm N}_{\rm eff}$ either thermally or non-thermally. Light Dirac neutrino models often lead to enhanced $\Delta{\rm N}_{\rm eff}$, some recent works on which can be found in Abazajian:2019oqj ; FileviezPerez:2019cyn ; Nanda:2019nqy ; Han:2020oet ; Luo:2020sho ; Borah:2020boy ; Adshead:2020ekg ; Luo:2020fdt ; Mahanta:2021plx ; Du:2021idh ; Biswas:2021kio ; Borah:2022obi ; Li:2022yna ; Das:2023yhv . Depending upon the strength of Yukawa couplings associated with $\nu_{R}$, we can have either thermal or non-thermal contribution to $\Delta{\rm N}_{\rm eff}$ as we discuss in the following subsections.
V.1 Thermal $\Delta N_{\rm eff}$
Figure 9: Interaction rate of $\nu_{R}$ compared to the Hubble expansion rate, indicating its decoupling profile for two sets of benchmark values.
We have three generations of right handed neutrinos ($\nu_{R_{i}}$) in our model, and their abundance in the Universe will be decided by the interaction strength and the masses of the particle with which they interact. The last three interaction terms of the Lagrangian (Eq. (II)) are the only three interactions that $\nu_{R}$ have with $N$ and $\chi$. The relative strength between the two types of Yukawa couplings $y_{i}$ and $y_{i}^{\prime}$ leads to three distinctive classes (Case-A, Case-B, Case-C) as we discussed before, in the context of neutrino mass and DM. With sufficiently large Yukawa coupling ($y_{i}^{\prime}$) as in Case-B, $\nu_{R}$ may enter the thermal bath with the SM degrees of freedom. As the Universe expands and the temperature falls, the scattering processes that kept $\nu_{R}$ in the thermal bath eventually go out of equilibrium, leading to the decoupling of $\nu_{R}$. For the combined scattering processes of $N_{L}\bar{\nu_{R}}\leftrightarrow N_{L}\bar{\nu_{R}}$ and $\chi\bar{\nu_{R}}\leftrightarrow\chi\bar{\nu_{R}}$, we check the decoupling or thermalization profile with two sets of benchmark parameters (masses and couplings) as shown in Fig. 9. As $N$ and $\chi$ masses lies at $\mathcal{O}$(TeV) range, it leads to the decoupling of $\nu_{R}$ from the thermal bath very early. Such early decoupling, above the electroweak scale, leads to a fixed value $\Delta N_{\rm eff}=0.14$ for three generations of $\nu_{R}$. Clearly, this remains within reach of several future CMB experiments mentioned earlier.
V.2 Non-thermal $\Delta N_{\rm eff}$
It is also possible to have observable $\Delta N_{\rm eff}$ when $\nu_{R}$ does not enter the thermal bath. In such a case, $\nu_{R}$ can be produced non-thermally from the decay of the scalar $\chi$ as $\chi\rightarrow N\nu_{R_{i}}$ or from decay of $N$ as $N\rightarrow\chi\nu_{R_{i}}$ depending upon our choice of fermion or scalar DM. Similar contributions can also arise from scatterings. With small Yukawa couplings for the $\bar{N_{L}}\chi\nu_{R}$ interaction (as in Case-A and Case-C), $\nu_{R}$ can not enter the thermal bath via the $N_{L}\bar{\nu_{R}}\leftrightarrow N_{L}\bar{\nu_{R}}$ (or $\chi\bar{\nu_{R}}\leftrightarrow\chi\bar{\nu_{R}}$) scattering process. In such a situation, where the scalar $\chi$ is in the thermal bath, $\nu_{R}$ are produced via the decay $\chi\rightarrow N\nu_{R}$ (assuming fermion DM scenario).
With $\chi$ in equilibrium and subsequently producing $N$ and $\nu_{R}$, the relevant Boltzmann equation for the comoving density of $\nu_{R}$ of flavour $\alpha$ is given by Biswas:2022vkq ,
$\frac{d\tilde{Y}_{\nu^{\alpha}_{R}}}{dx}=\frac{\beta}{x\mathcal{H}s^{1/3}}\langle E\Gamma\rangle Y^{\rm eq}_{\chi}.$ (15)
Here, $x=\frac{M_{\chi}}{T}$, $\mathcal{H}$ is the Hubble parameter in the radiation dominated era and $\beta$ is defined as $\beta=\big{[}1+T\frac{dg_{s}/dT}{3g_{s}}\big{]}$ with $g_{s}$ being the effective relativistic entropy degrees of freedom. For the sub-eV scale $\nu_{R}$, which remains relativistic till recombination, the comoving energy density is defined as $\tilde{Y}_{\nu^{\alpha}_{R}}=\rho_{\nu^{\alpha}_{R}}/s^{4/3}$ with $\rho_{\nu^{\alpha}_{R}},s$ denoting energy density of $\nu^{\alpha}_{R}$ and entropy density of the Universe respectively. On the right hand side of Eq. (15), $Y^{\rm eq}_{\chi}\equiv n^{\rm eq}_{\chi}/s$ denotes the equilibrium comoving density of $\chi$. The quantity $\langle E\Gamma\rangle$ is defined as
$\langle E\Gamma\rangle=\frac{M_{\chi}}{2}\left(1-\frac{M^{2}_{N}}{M^{2}_{\chi}}\right)\Gamma_{\chi}$ (16)
where $\Gamma_{\chi}$ denotes the partial decay width for the process $\chi\rightarrow N\nu_{R}$. The $\nu_{R}$ produced from $\chi$ decay will contribute to this additional radiation density. Therefore, with $\rho_{R}=\rho_{\rm SM}+\rho_{\nu_{R}}$, $\Delta N_{\rm eff}$ for one species of $\nu_{R}$ (indicated by flavour index $\alpha$) calculated at the CMB epoch $(T_{\rm CMB})$ is
$\Delta N^{\alpha}_{\rm eff}$ $=$ $N^{\alpha}_{\rm eff}-N_{\rm eff}^{\rm SM}$ (17)
$=$ $2\times\Big{(}\frac{\rho_{\nu^{\alpha}_{R}}}{\rho_{\nu_{L}}}\Big{)}_{T_{\rm CMB}}$
$=$ $2\times\Big{(}\frac{\rho_{\nu^{\alpha}_{R}}}{\rho_{\nu_{L}}}\Big{)}_{T_{\rm BBN}}$
$=$ $2\times\Big{(}\frac{s^{4/3}\tilde{Y}_{\nu^{\alpha}_{R}}}{\rho_{\nu_{L}}}\Big{)}_{T_{\rm BBN}}$
where, the extrapolation from the BBN epoch $(T_{\rm BBN})$ to CMB epoch $(T_{\rm CMB})$ is justified by the fact that both $\rho_{\nu_{L}},\rho_{\nu_{R}}$ redshift similarly for $T<T_{\rm BBN}$. One can simply calculate the energy density ratio at BBN epoch $\sim\mathcal{O}({10})$ MeV as $\nu_{R}$ production stops much earlier. The numerical factor 2 in Eq. (17) indicate the degrees of freedom of each $\nu_{R}$. The net contribution to $\Delta N_{\rm eff}$ can be obtained by summing over flavour index $\alpha$ as $\Delta N_{\rm eff}=\sum_{\alpha}\Delta N^{\alpha}_{\rm eff}$. Ignoring the temperature dependence of the relativistic degrees of freedom, one can solve Eq. (15) analytically to find $\Delta N_{\rm eff}$ Biswas:2022vkq .
Figure 10: Contribution to non-thermal $\Delta N_{\rm eff}$ for the Case-C where $y_{i}^{\prime}\sim\mathcal{O}(10^{-6})$, considering the decay $\chi\rightarrow\nu_{R}N$.
The decoupling profile for $\nu_{R}$ for two different benchmark choices of Yukawa coupling $y_{i}^{\prime}$ is shown in Fig. 9. For scalar ($\chi$) and fermion ($N$) masses in the TeV ballpark, we find that for $y_{i}^{\prime}\leq\mathcal{O}(10^{-3})$, the interaction $N_{L}\bar{\nu_{R}}\leftrightarrow N_{L}\bar{\nu_{R}}$ can not attain thermal equilibrium in the early Universe. Therefore, Case-A with $y_{i}^{\prime}\sim\mathcal{O}(10^{-11})$, $\nu_{R}$ will never enter the thermal bath, hence, thermal $\Delta N_{\rm eff}$ is not possible. On the other hand, $\nu_{R}$ produced non-thermally from the decay of either $\chi$ or $N$ gives rise to negligible contribution to $\Delta N_{\rm eff}$. As mentioned before, the Case-B, with $0.01\leq y_{1}^{\prime}\leq 0.1$, $0.03\leq y_{2}^{\prime}\leq 0.1$, $0.03\leq y_{3}^{\prime}\leq 0.1$, we have thermal contribution to $\Delta N_{\rm eff}$. Due to the decoupling of $\nu_{R}$ much above the electroweak scale, we have a fixed contribution $\Delta N_{\rm eff}=0.14$ for Case-B. Finally, in Case-C, Yukawa couplings are of the size $\mathcal{O}(10^{-6})$, and the corresponding interactions do not allow $\nu_{R}$ to enter the thermal bath. Considering the decay of $\chi\rightarrow\nu_{R}N$, we have shown the non-thermal contributions to $\Delta N_{\rm eff}$ in Fig. 10. Clearly, some part of the parameter space is already ruled out by PLANCK 2018 and DESI 2024 limits while most of the remaining parameter space remains within reach of future CMB experiments like SPT-3G, CMB-S4 and CMB-HD. As there are multiple $\mathcal{Z}_{2}$-odd scalars, it is possible to have $\chi_{i}\rightarrow\nu_{R}N$ type decays from one of the heavier scalars irrespective of fermion or scalar DM. While contribution to $\Delta N_{\rm eff}$ from scalar decay can be calculated simply by considering the scalar $\chi_{i}$ to be in equilibrium, contributions from the decay $N\rightarrow\nu_{R}\chi_{j}$ will be more involved as the number density of non-thermal $N$ needs to be found by solving the respective Boltzmann equation Biswas:2022vkq .
To summarise, observable $\Delta N_{\rm eff}$ is possible in Case-B and Case-C. However, fermion DM is not possible in these two cases irrespective of freeze-out or freeze-in mechanism. On the other hand, fermion DM produced via freeze-out mechanism is viable in Case-A even though the $\Delta N_{\rm eff}$ remains out of reach from future CMB experiments. This conflict between observable $\Delta N_{\rm eff}$ and DM relic can be avoided if we consider scalar DM. Scalar DM allows light scalars in sub-TeV ballpark with observable direct detection rates while keeping collider prospects more promising compared to fermion DM scenario where all $\mathcal{Z}_{2}$-odd particles remain in multi-TeV ballpark.
VI Other detection prospects
Figure 11: One-loop contribution to charged lepton flavour violating $\ell_{\eta}\rightarrow\ell_{\alpha}\gamma$ process
In addition to the detection possibilities of DM and $\Delta N_{\rm eff}$, the model can have other detection aspects too. We briefly comment on such possibilities here.
Firstly, we can have a lepton flavour violating (LFV) process at one-loop level, as shown in Fig. 11. The corresponding decay width can be calculated using the prescriptions given in Kuno:1999jp ; Guo:2020qin . With relevant Yukawa couplings of $y_{i}\sim\mathcal{O}(10^{-1}-10^{-2})$, as in Case-A, it gives the branching ratio $\text{Br}(\mu\rightarrow e\gamma)\sim\mathcal{O}(10^{-15})$, which is close to the recent upper bound indicated by MEG-II experiment $\text{Br}(\mu\rightarrow e\gamma)<3.1\times 10^{-13}$ MEGII:2023ltw . For Case-B and Case-C with smaller Yukawa couplings $y_{i}\sim\mathcal{O}(10^{-11})$ and $y_{i}\sim\mathcal{O}(10^{-7})$, we get very small branching ratio $\text{Br}(\mu\rightarrow e\gamma)$ as $\mathcal{O}(10^{-57})$ and $\mathcal{O}(10^{-41})$ respectively, beyond experimental reach in near future.
Due to the presence of several new scalars, admixtures of singlet and doublet, we can have different collider prospects as well. If scalar DM mass $(M_{\chi})$ is lighter than the SM Higgs mass $(M_{h})$ such that $h\rightarrow\chi\chi$ decay is allowed, it can contribute to invisible Higgs decay. Such invisible Higgs decay is currently constrained at Br$(h\rightarrow{\rm invisible})<0.14$ ATLAS:2022yvh . Production of $\mathcal{Z}_{2}$-odd scalars in colliders like the large hadron collider (LHC) can give rise to pure leptonic final states plus missing transverse energy (MET) Gustafsson:2012aj ; Datta:2016nfz , hadronic final states plus MET Poulose:2016lvz or a mixture of both. Such MET can arise from light Dirac neutrinos or scalar/fermion DM in our scenario. It is also possible to have more exotic signatures like tri-lepton plus MET Hashemi:2016wup and mono-jet Belyaev:2016lok ; Belyaev:2018ext . It should be noted that we also have new scalars in the model which are $\mathcal{Z}_{2}$-even. They can lead to additional final states, similar to generic two Higgs doublet models. The new charged scalars can also modify the diphoton decay width of the SM Higgs, currently constrained at $\frac{{\rm BR}(h\rightarrow\gamma\gamma)_{\rm expt}}{{\rm BR}(h\rightarrow\gamma\gamma)_{\rm SM}}=1.12\pm 0.09$ CMS:2021kom .
Another interesting detection prospect arises in the gravitational wave (GW) frontier. Spontaneous breaking of discrete symmetries like $A_{4}\times Z_{2}$, as we have in our model, can lead to formation of topological defects like domain walls (DW). While DW, if stable or allowed to dominate the energy density of the Universe, will be in conflict with cosmological observations, they can be made to disappear by introducing small bias, in the form of explicit symmetry breaking terms. Such annihilating or collapsing DW then generates stochastic GW which can be probed at ongoing or future experiments. GW signatures of minimal Dirac seesaw model was studied in earlier works Barman:2022yos ; Barman:2023fad while similar studies in the context of flavour symmetric models like $A_{4}$ can be found in Gelmini:2020bqg . The bias or explicit symmetry breaking terms introduced for DW disappearance may also lead to interesting indirect detection signatures of DM due to its decay into different standard model particles [^2] See King:2023ayw ; Borah:2024kfn for related discussions.. While the model has several detection prospects, it can also be falsified by future observation of neutrinoless double beta decay.
Yukawa Neurino Observable $\Delta N_{\rm eff}$ Observable $\Delta N_{\rm eff}$ Fermion Scalar Observable
Case Coupling Data (Thermal) (Non-thermal) DM DM LFV
A $y_{i}\gg y_{i}^{\prime}$ Yes No No Yes Yes Yes
B $y_{i}\ll y_{i}^{\prime}$ Yes Yes Not applicable No Yes No
C $y_{i}\simeq y_{i}^{\prime}$ Yes No Yes No Yes No
Table 2: Summary of results for three benchmark scenarios.
VII Conclusion
We have proposed a new model of light Dirac neutrinos based on non-Abelian discrete flavour symmetry $A_{4}$ augmented with $Z_{2}\times U(1)_{L}$ where the symmetry breaking pattern leaves a remnant $\mathcal{Z}_{2}\subset A_{4}$ symmetry responsible for stabilising DM. Using the minimal flavon and additional fermion content required for realising neutrino data, we take both tree level and one-loop contributions to light Dirac neutrino masses. While tree level or type-I seesaw leads to a rank-1 Dirac neutrino mass matrix with only one non-zero mass eigenvalue, scotogenic or one-loop contributions can generate the other masses. Our proposal, therefore, provides the first generalisation of discrete dark matter and scoto-seesaw scenarios to light Dirac neutrinos.
The remnant $\mathcal{Z}_{2}$ symmetry allows either fermion or scalar DM candidates. While scalar DM is dominantly produced thermally due to gauge and scalar portal interactions, fermion DM production relies primarily on Yukawa portal interactions. The same Yukawa portal interactions can also lead to thermal or non-thermal production of right handed neutrinos, which contribute to the effective relativistic degrees of freedom $\Delta N_{\rm eff}$. To illustrate this, we consider three different benchmark scenarios of such Yukawa couplings and perform a detailed study of neutrino mass, mixing together with DM and $\Delta N_{\rm eff}$. We also check the scope of other detection prospects like LFV decays. Table 2 summarises our results. While one of the sub-classes of Yukawa couplings, namely Case-A, allows fermion singlet DM with observable LFV rates, the corresponding $\Delta N_{\rm eff}$ remains negligible. The other two sub-classes allow scalar DM only but with observable $\Delta N_{\rm eff}$ at future CMB experiments. The LFV rates however remain tiny in these two sub-classes. Due to the rich particle spectrum in terms of several scalar doublets and singlets, the model can also have rich collider phenomenology. On the other hand, spontaneous breaking of discrete symmetries can have interesting consequences in cosmology due to the formation of domain walls. Such walls, which disappear due to soft-breaking terms, can emit stochastic gravitational waves which can be probed at different ongoing and future experiments. The same soft-breaking terms can also lead to decay of DM, keeping indirect detection prospects alive. We leave a comprehensive analysis of such additional detection prospects of our model to future works.
Acknowledgements.
The work of D.B. is supported by the Science and Engineering Research Board (SERB), Government of India grants MTR/2022/000575 and CRG/2022/000603. The work of B.K. has been supported in part by the Polish National Science Center (NCN) under grant 2020/37/B/ST2/02371, the Freedom of Research, Mobility of Science, Research Excellence Initiative of the University of Silesia in Katowice and EU COST Action (CA21136). S.M. acknowledges the financial support from the National Research Foundation (NRF) grant funded by the Korea government (MEST) NRF-2022R1A2C1005050.
Appendix A $A_{4}$ Multiplication rules
All finite groups are completely characterized by means of a set of elements called generators of the group and a set of relations, so that all the elements of the group are given as product of the generators. The group $A_{4}$ consists of the even permutations of four objects and then contains 4!/2 = 12 elements. The generators are S and T with the relations $S^{2}=T^{3}=(ST)^{3}=\mathcal{I}$, then the elements are $1,S,T,ST,TS,T^{2},ST^{2},STS,TST,T^{2}S,TST^{2},T^{2}ST$.
It has four irreducible representations: three one-dimensional and one three-dimensional which are denoted by $\bf{1},\bf{1^{\prime}},\bf{1^{\prime\prime}}$ and $\bf{3}$ respectively. Under $A_{4}$, one-dimensional unitary representations are given as,
$1\displaystyle 1\quad S=1\quad T=$ 1
$1^{\prime}\quad S=1\quad T=\omega$
$1^{\prime\prime}\quad S=1\quad T=\omega^{2}$
$\omega=e^{i\frac{2\pi}{3}}$ is cube root of unity. In the basis, where $S$ is a real diagonal,
$S=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&-1\end{pmatrix}\quad T=\begin{pmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{pmatrix}$ (18)
The multiplication rules of the irreducible representations are given by Altarelli:2010gt
$\bf{1}\otimes\bf{1}=\bf{1},\bf{1^{\prime}}\otimes\bf{1^{\prime}}=\bf{1^{\prime\prime}},\bf{1^{\prime}}\otimes\bf{1^{\prime\prime}}=\bf{1},\bf{1^{\prime\prime}}\otimes\bf{1^{\prime\prime}}=\bf{1^{\prime}},\bf{3}\otimes\bf{3}=\bf{1}+\bf{1^{\prime}}+\bf{1^{\prime\prime}}+\bf{3}_{1}+\bf{3}_{2}$ (19)
where ${\bf a}$ and ${\bf s}$ in the subscript corresponds to anti-symmetric and symmetric parts respectively. Now, if we have two triplets as $A=(a_{1},a_{2},a_{3})^{T}$ and $B=(b_{1},b_{2},b_{3})^{T}$ respectively, their direct product can be decomposed into the direct sum mentioned above. The product rule for this two triplets in the $S$ diagonal basis [^3] Here $S$ is a $3\times 3$ diagonal generator of $A_{4}$. can be written as He:2006dk ; Memenga:2013vc ; Ishimori:2010au
[…]
where $\omega$ ($=e^{2i\pi/3}$) is the cube root of unity.
Appendix B The scalar potential
The tree level scalar potential of our model can be written as
$V(H,\eta,\chi)$ $=$ $-\mu_{h}^{2}(H^{\dagger}H)-\mu_{\eta}^{2}(\eta^{\dagger}\eta)_{1}-\mu_{\chi}^{2}(\chi\chi)_{1}+\lambda_{H}(H^{\dagger}H)^{2}+\lambda_{\eta 1}(\eta^{\dagger}\eta)_{1}(\eta^{\dagger}\eta)_{1}+\lambda_{\eta 2}(\eta^{\dagger}\eta)_{1^{\prime}}(\eta^{\dagger}\eta)_{1^{\prime\prime}}$ (25)
$+\lambda_{\eta 3}(\eta^{\dagger}\eta)_{3_{1}}(\eta^{\dagger}\eta)_{3_{1}}+\lambda_{\eta 4}[(\eta^{\dagger}\eta)_{3_{1}}(\eta^{\dagger}\eta)_{3_{2}}+{\rm h.c.}]+\lambda_{\eta 5}[(\eta^{\dagger}\eta)_{3_{2}}(\eta^{\dagger}\eta)_{3_{2}}+{\rm h.c.}]$
$+\lambda_{\eta H1}(\eta^{\dagger}\eta)_{1}(H^{\dagger}H)+\lambda_{\eta H2}(\eta^{\dagger}H)(H^{\dagger}\eta)+\lambda_{\eta H3}[(\eta^{\dagger}H)(\eta^{\dagger}H)+{\rm h.c.}]$
$+\lambda_{\eta H4}\big{[}(\eta^{\dagger}\eta)_{3_{1}}H^{\dagger}\eta+{\rm h.c.}\big{]}+\lambda_{\eta H5}\big{[}(\eta^{\dagger}\eta)_{3_{2}}H^{\dagger}\eta+{\rm h.c.}\big{]}+\lambda_{\chi 1}(\chi\chi)_{1}(\chi\chi)_{1}$
$+\lambda_{\chi 2}(\chi\chi)_{1^{\prime}}(\chi\chi)_{1^{\prime\prime}}+\lambda_{\chi 3}(\chi\chi)_{3_{1}}(\chi\chi)_{3_{1}}+\lambda_{\chi H1}(\chi\chi)_{1}(H^{\dagger}H)+\lambda_{\eta\chi 1}(\eta^{\dagger}\eta)_{1}(\chi\chi)_{1}$
$+\lambda_{\eta\chi 2}\big{[}(\eta^{\dagger}\eta)_{1^{\prime}}(\chi\chi)_{1^{\prime\prime}}+{\rm h.c.}\big{]}+\lambda_{\eta\chi 3}\big{[}(\eta^{\dagger}\eta)_{1^{\prime\prime}}(\chi\chi)_{1^{\prime}}+{\rm h.c.}\big{]}+\lambda_{\eta\chi 4}(\eta^{\dagger}\eta)_{3_{1}}(\chi\chi)_{3_{1}}$
$+\lambda_{\eta\chi 5}\big{[}(\eta^{\dagger}\eta)_{3_{2}}(\chi\chi)_{3_{1}}+{\rm h.c.}\big{]}+\lambda_{\eta H\chi}\big{[}(\chi\chi)_{3_{1}}(H^{\dagger}\eta)+{\rm h.c.}\big{]}$
where, the product of two triplets contracted into one of the two triplet representations of $A_{4}$ is indicated by $[...]_{3_{1,2}}$, and the product of two triplets contracted into a singlet is represented by $[...]_{1,1^{\prime},1^{\prime\prime}}$. Among the $A_{4}$ triplet scalars $\eta=(\eta_{1},\eta_{2},\eta_{3})$ and $\chi=(\chi_{1},\chi_{2},\chi_{3})$, only $\eta_{1}$ and $\chi_{1}$ are getting VEVs $v_{\eta}$ and $v_{\chi}$ respectively. The scalar fields can be can be expressed as
$H$ $=$ $\begin{pmatrix}h^{\pm}\\ (v+h+iA)/\sqrt{2}\end{pmatrix};$ (26)
$\eta_{1}$ $=$ $\begin{pmatrix}\eta_{1}^{\pm}\\ (v_{\eta}+h_{1}+iA_{1})/\sqrt{2}\end{pmatrix},\quad\eta_{2,3}=\begin{pmatrix}\eta_{2,3}^{\pm}\\ (h_{2,3}+iA_{2,3})/\sqrt{2}\end{pmatrix};$ (27)
$\chi_{1}$ $=$ $(v_{\chi}+x_{1})/\sqrt{2},\quad\chi_{2,3}=x_{2,3}/\sqrt{2}.$ (28)
The minimisation conditions for the scalars obtaining non-zero VEVs can be written as
$\mu^{2}_{H}=$ $\lambda_{H}v^{2}+(\lambda_{\eta H1}+\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v^{2}_{\eta}}{2}+\lambda_{\chi H1}\frac{v^{2}_{\chi}}{2},$
$\mu^{2}_{\eta}=$ $(\lambda_{\eta 1}+\lambda_{\eta 2})v^{2}_{\eta}+(\lambda_{\eta H1}+\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v^{2}}{2}+(\lambda_{\eta\chi 1}+2(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3}))\frac{v^{2}_{\chi}}{2},$ (29)
$\mu^{2}_{\chi}=$ $(\lambda_{\chi 1}+\lambda_{\chi 2})v^{2}_{\chi}+\lambda_{\chi H1}\frac{v^{2}}{2}+(\lambda_{\eta\chi 1}+2(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3}))\frac{v^{2}_{\eta}}{2}.$
There are seven neutral scalars $(h,h_{1},h_{2},h_{3},x_{1},x_{2},x_{3})$ out of which four are $\mathcal{Z}_{2}$-odd while the remaining three are $\mathcal{Z}_{2}$-even. Similarly, there is one physical pseudo-scalar out of the $\mathcal{Z}_{2}$-even combination of $(A,A_{1})$ while $(A_{2},A_{3})$ constitute $\mathcal{Z}_{2}$-odd pseudoscalars. The $\mathcal{Z}_{2}$-even scalar mass matrix squared in the $(h,h_{1},x_{1})$ basis can be written as
$\mathcal{M}^{2}_{S+}=\begin{pmatrix}M^{2}_{hh}&M^{2}_{hh_{1}}&M^{2}_{hx_{1}}\\ M^{2}_{hh_{1}}&M^{2}_{h_{1}h_{1}}&M^{2}_{h_{1}x_{1}}\\ M^{2}_{hx_{1}}&M^{2}_{h_{1}x_{1}}&M^{2}_{x_{1}x_{1}}\\ \end{pmatrix}_{3\times 3}$ (30)
where,
$M^{2}_{hh}=2\lambda_{H}v^{2},\,\,M^{2}_{hh_{1}}=(\lambda_{\eta H1}+\lambda_{\eta H2}+2\lambda_{\eta H3})v_{\eta}v,\,\,M^{2}_{hx_{1}}=\lambda_{\chi H1}v_{\chi}v,$
$M^{2}_{h_{1}h_{1}}=2(\lambda_{\eta 1}+\lambda_{\eta 2})v^{2}_{\eta},\,\,M^{2}_{h_{1}x_{1}}=(\lambda_{\eta\chi 1}+2(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3}))v_{\chi}v_{\eta},\,\,M^{2}_{x_{1}x_{1}}=2(\lambda_{\chi 1}+\lambda_{\chi 2})v^{2}_{\chi}.$
This mass squared matrix can be diagonalised by $3\times 3$ orthogonal matrix involving three angles. Similarly, the $\mathcal{Z}_{2}$-odd scalar mass matrix squared in the $(h_{2},h_{3},x_{2},x_{3})$ basis can be written as
$\mathcal{M}^{2}_{S-}=\begin{pmatrix}M^{2}_{h_{2}h_{2}}&M^{2}_{h_{2}h_{3}}&M^{2}_{h_{2}x_{2}}&M^{2}_{h_{2}x_{3}}\\ M^{2}_{h_{2}h_{3}}&M^{2}_{h_{3}h_{3}}&M^{2}_{h_{3}x_{2}}&M^{2}_{h_{3}x_{3}}\\ M^{2}_{h_{2}x_{2}}&M^{2}_{h_{3}x_{2}}&M^{2}_{x_{2}x_{2}}&M^{2}_{x_{2}x_{3}}\\ M^{2}_{h_{2}x_{3}}&M^{2}_{h_{3}x_{3}}&M^{2}_{x_{2}x_{3}}&M^{2}_{x_{3}x_{3}}\\ \end{pmatrix}_{4\times 4},$ (31)
where
$M^{2}_{h_{2}h_{2}}=$ $M_{h_{3}h_{3}}^{2}=\frac{1}{2}((-3\lambda_{\eta 2}+4\lambda_{\eta 3})v^{2}_{\eta}-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v_{\chi}^{2}),$
$M^{2}_{x_{2}x_{2}}=$ $M_{x_{3}x_{3}}^{2}=\frac{1}{2}[-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v_{\eta}^{2}+(-3\lambda_{\chi 2}+4\lambda_{\chi 3})v_{\chi}^{2}],$
$M^{2}_{h_{2}h_{3}}=$ $3\lambda_{\eta H4}vv_{\eta}$
$M^{2}_{h_{2}x_{2}}=$ $M_{h_{3}x_{3}}^{2}=\lambda_{\eta H\chi}vv_{\chi}+\lambda_{\eta\chi 4}v_{\eta}v_{\chi},$
$M_{h_{2}x_{3}}^{2}=$ $M_{h_{3}x_{2}}^{2}=\lambda_{\eta H\chi}vv_{\chi}$
$M_{x_{2}x_{3}}^{2}=$ $\lambda_{\eta H\chi}vv_{\eta}$
This mass squared matrix can similarly be diagonalised by a $4\times 4$ orthogonal matrix involving six angles, in general, leading to physical states $\zeta_{i}$. In the scalar DM case, the lightest eigenstate of the $4\times 4$ mass matrix given above becomes the DM candidate. Similar mass matrices can be written for both $\mathcal{Z}_{2}$-odd and $\mathcal{Z}_{2}$-even pseudoscalars as well.
The mass matrix squared for pseduo-scalars in the ($A,A_{1},A_{2},A_{3}$) basis can be written as
$\mathcal{M}_{PS}^{2}=\begin{pmatrix}M^{2}_{AA}&M^{2}_{AA_{1}}&0&0\\ M^{2}_{AA_{1}}&M^{2}_{A_{1}A_{1}}&0&0\\ 0&0&M^{2}_{A_{2}A_{2}}&M^{2}_{A_{2}A_{3}}\\ 0&0&M^{2}_{A_{2}A_{3}}&M^{2}_{A_{3}A_{3}}\end{pmatrix}_{4\times 4}$ (32)
where
$M^{2}_{AA}=$ $-2\lambda_{\eta H3}v_{\eta}^{2},\quad M^{2}_{A_{1}A_{1}}=-2\lambda_{\eta H3}v^{2},\quad M^{2}_{AA_{1}}=2\lambda_{\eta H3}vv_{\eta},$
$M_{A_{2}A_{2}}^{2}=$ $M_{A_{3}A_{3}}^{2}=\frac{1}{2}[-4\lambda_{\eta H3}v^{2}-(3\lambda_{\eta 2}+8\lambda_{\eta 5})v_{\eta}^{2}-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v_{\chi}^{2}],$
$M^{2}_{A_{2}A_{3}}=$ $\lambda_{\eta H4}vv_{\eta}.$
Similarly for the charged scalars, the mass matrix squared in the ($h^{\pm},\eta_{1}^{\pm},\eta_{2}^{\pm},\eta_{3}^{\pm}$) basis can be expressed as
$\mathcal{M}^{2}_{CS}=\begin{pmatrix}M^{2}_{h^{\pm}h^{\mp}}&M^{2}_{h^{\pm}\eta_{1}^{\mp}}&0&0\\ M^{2}_{h^{\pm}\eta_{1}^{\mp}}&M^{2}_{\eta_{1}^{\pm}\eta_{1}^{\mp}}&0&0\\ 0&0&M^{2}_{\eta_{2}^{\pm}\eta_{2}^{\mp}}&M^{2}_{\eta_{2}^{\pm}\eta_{3}^{\mp}}\\ 0&0&M^{2}_{\eta_{2}^{\pm}\eta_{3}^{\mp}}&M^{2}_{\eta_{3}^{\pm}\eta_{3}^{\mp}}\end{pmatrix}_{4\times 4}$ (33)
$M^{2}_{h^{\pm}h^{\mp}}=$ $-(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v_{\eta}^{2}}{2},\quad M^{2}_{\eta_{1}^{\pm}\eta_{1}^{\mp}}=-(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v^{2}}{2},$
$M_{\eta_{2}^{\pm}\eta_{2}^{\mp}}^{2}=$ $M^{2}_{\eta_{3}^{\pm}\eta_{3}^{\mp}}=-(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v^{2}}{2}-3\lambda_{\eta 2}\frac{v_{\eta}^{2}}{2}-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})\frac{v_{\chi}^{2}}{2},$
$M^{2}_{h^{\pm}\eta_{1}^{\mp}}=$ $(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{vv_{\eta}}{2},\quad M^{2}_{\eta_{2}^{\pm}\eta_{3}^{\mp}}=\lambda_{\eta H4}vv_{\eta}.$
Clearly, the $\mathcal{Z}_{2}$-even $2\times 2$ blocks of $\mathcal{M}^{2}_{PS}$ and $\mathcal{M}^{2}_{CS}$ have rank 1 indicating the massless Goldstone modes.
Appendix C Neutrino mass matrix
Combining the tree level and one-loop contributions, the complete neutrino mass matrix can be written as
$m_{\nu}$ $=$ $m_{\nu 0}+m^{\rm loop}_{\nu}$ (34)
$=$ $\begin{pmatrix}(m_{\nu})_{11}&(m_{\nu})_{12}&(m_{\nu})_{13}\\ (m_{\nu})_{21}&(m_{\nu})_{22}&(m_{\nu})_{23}\\ (m_{\nu})_{31}&(m_{\nu})_{32}&(m_{\nu})_{33}\end{pmatrix}$ (35)
where,
$(m_{\nu})_{11}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{1}y^{\prime}_{1}+y_{1}y^{\prime}_{1}M_{N}(\mathcal{F}_{1}+\mathcal{F}_{2}+\mathcal{F}_{3}),$ (36)
$(m_{\nu})_{12}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{1}y^{\prime}_{2}+y_{1}y^{\prime}_{2}M_{N}(\mathcal{F}_{1}+\omega\mathcal{F}_{2}+\omega^{2}\mathcal{F}_{3}),$ (37)
$(m_{\nu})_{13}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{1}y^{\prime}_{3}+y_{1}y^{\prime}_{3}M_{N}(\mathcal{F}_{1}+\omega^{2}\mathcal{F}_{2}+\omega\mathcal{F}_{3}),$ (38)
$(m_{\nu})_{21}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{2}y^{\prime}_{1}+y_{2}y^{\prime}_{1}M_{N}(\mathcal{F}_{1}+\omega\mathcal{F}_{2}+\omega^{2}\mathcal{F}_{3}),$ (39)
$(m_{\nu})_{22}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{2}y^{\prime}_{2}+y_{2}y^{\prime}_{2}M_{N}(\mathcal{F}_{1}+\omega^{2}\mathcal{F}_{2}+\omega\mathcal{F}_{3}),$ (40)
$(m_{\nu})_{23}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{2}y^{\prime}_{3}+y_{2}y^{\prime}_{3}M_{N}(\mathcal{F}_{1}+\mathcal{F}_{2}+\mathcal{F}_{3}),$ (41)
$(m_{\nu})_{31}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{3}y^{\prime}_{1}+y_{3}y^{\prime}_{1}M_{N}(\mathcal{F}_{1}+\omega^{2}\mathcal{F}_{2}+\omega\mathcal{F}_{3}),$ (42)
$(m_{\nu})_{32}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{3}y^{\prime}_{2}+y_{3}y^{\prime}_{2}M_{N}(\mathcal{F}_{1}+\mathcal{F}_{2}+\mathcal{F}_{3}),$ (43)
$(m_{\nu})_{33}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{3}y^{\prime}_{3}+y_{3}y^{\prime}_{3}M_{N}(\mathcal{F}_{1}+\omega\mathcal{F}_{2}+\omega^{2}\mathcal{F}_{3}).$ (44)
Now, substituting $y_{2}=y_{3}$ and $y^{\prime}_{2}=y^{\prime}_{3}$ in Eq. (36) - Eq. (44) and assuming all Yukawa couplings to be real, we find the following relations:
$(m_{\nu})_{11}=(m_{\nu})^{*}_{11},~{}(m_{\nu})_{12}=(m_{\nu})^{*}_{13},~{}(m_{\nu})_{21}=(m_{\nu})^{*}_{31},$
$(m_{\nu})_{23}=(m_{\nu})^{*}_{32},(m_{\nu})_{22}=(m_{\nu})^{*}_{33}.$ (45)
These conditions give rise to a $\mu-\tau$ reflection symmetric mass matrix for Dirac neutrinos with maximal atmospheric mixing angle and Dirac CP phase.